Use the following information for questions 1-2:

The following data from the just completed year are taken from the accounting records of Mason Company.

Please use the information below:

Sales $524,000

Direct Labor Cost $70,000

Raw Materials Purchased $118,000

Selling Expense $140,000

Administrative Expense $63,000

Manufacturing Overhead Applied to Work in Process $90,000

Actual Manufacturing Overhead Cost $100,000

Inventories Beginning of Year End of Year

Raw Materials $7,000 $15,000

Work in Process ? $5,000

Finished Goods $20,000 ?

Additional information: For the year, the cost of goods manufactured totaled $280,000; the unadjusted cost of goods sold totaled $260,000.

The company’s overapplied or underapplied overhead is closed entirely to Cost of Goods Sold.

1. How much was the Company’s Work in Process inventory balance at the beginning of the year?

A. $2,000 B. $10,000 C. $15,000 D. $20,000 E. None of the above

2. How much was the Company’s adjusted Cost of Goods Sold at the end of the year?

A. $280,000 B. $240,000 C. $250,000 D. $270,000 E. None of the above

3. Botox Inc. had the following financial results for 2013:

Sales Revenue $1,680,000

Contribution Margin $1,008,000

Net Operating Profit $108,000

What is Botox’s margin of safety in dollars?

A. $1,800,000 B. $360,000 C. $180,000 D. $240,000 E. $720,000

4. A company’s total overhead cost at various levels of activity is presented below:

Month Machine-Hours Total Overhead Cost

March ...................................... 5,000 $21,750

April ......................................... 4,000 $19,000

May ......................................... 6,000 $25,000

June ........................................ 7,000 $28,750

Assume that the overhead cost above consists of utilities, supervisory salaries, and maintenance.

The breakdown of these costs at the 4,000 machine-hour level of activity is as follows:

Utilities (variable) ............................... $ 8,000

Supervisory salaries (fixed) ................. 4,000

Maintenance (mixed) .......................... 7,000

Total overhead cost ............................ $19,000

Suppose the company uses the high-low method to estimate a cost formula for maintenance.

What is the maintenance cost the company expects to incur at an activity level of 5,500 machine hours?

A. $12,875 B. $19,875 C. $23,875 D. $8,875 E. None of the above

Could you please be kind enough and show me the solutions and answers in detail? Thank you so much for reading this and have a good one!
Answer #1

As per policy only one question is allowed to answer at a time, so answering 1-3 :

1) Beginning WIP Inventory balance=Cogm + c WIP + c Mat. - Mat purch - o Mat - D Lab - O/H applied
280000+5000+15000-118000-7000-70000-90000=(C) $15000
2) Adjusted Cost of Goods sold = Unadjusted COGS -(Actual manufacturing overhead cost - Manufacturing overhead applied
260000 - (100000 - 90000) = (C) $250000
3) Margin of safety = Sales revenue - BE sales
BE sales = FC / contribution margin = (conti $1008000 - NOP $108000)/(1008000/1680000)=1500000
Margin of safety = 1680000 - 1500000 = (C) $180000